# Raoyuan Zhao MaiNLP, Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML)

### v2

@@ Appendix A Computational Resources and AI Assistance
We generate the texts with temperature $=0.7$, top-$p=0.9$, and top-$k=10$ for all LLMs that run locally. All models were executed on NVIDIA A100 GPUs, except for the LLaMA3-70B model, which was run on an H200 GPU. The DeepSeek and ChatGPT models were accessed via their respective public APIs.
@@ C.1 Performance of Entity Extraction
We randomly sampled 50 sentences for each of six topics across four languages: English, Chinese, Arabic, and Hindi, to manually evaluate the performance of ChatGPT-4o-mini on entity extraction. Each language was annotated by one native speaker of that language. All annotators were native speakers with a graduate-level education background. While only one annotator was involved per language, we ensured careful guidelines and consistency checks to support evaluation reliability. As shown in Table 8 below, the F1 scores across all languages are consistently high, demonstrating the reliability of entity extraction process.
@@ C.2 Performance of Wikidata Search
We computed the proportion of extracted entities that could not be matched to any valid QID. As shown in Table 9, the average proportion of missing QIDs ranges from approximately 26% to 35% across different topics. We manually reviewed the entities that could not be matched to any valid QID and found clear signs of hallucination. In particular, the model frequently generated non-existent items, such as fictional book titles, songs, or dishes. This phenomenon poses a challenge for future efforts to manually complete missing QIDs, as hallucinated entities cannot be trivially verified or linked to real-world knowledge bases.
@@ Appendix G Results of Culture Consensus Results
In Figure 2530, we present the average consensus scores across all mentioned countries and six topics. To aid interpretation, we use yellow color coding to highlight major Europe region. Notably, European languages consistently exhibit higher cultural consensus across models, suggesting stronger alignment in entity representations within this region.

### v3

@@ Appendix A Computational Resources and AI Assistance
We generate the texts with temperature =0.7=0.7, top-$p=0.9$, and top-$k=10$ for all LLMs that run locally. All models were executed on NVIDIA A100 GPUs, except for the LLaMA3-70B model, which was run on an H200 GPU. The DeepSeek and ChatGPT models were accessed via their respective public APIs.
@@ C.1 Performance of Entity Extraction
We randomly sampled 50 sentences for each of six topics across four languages: English, Chinese, Arabic, and Hindi, to manually evaluate the performance of ChatGPT-4o-mini on entity extraction. Each language was annotated by one native speaker of that language. All annotators were native speakers with a graduate-level education background. While only one annotator was involved per language, we ensured careful guidelines and consistency checks to support evaluation reliability. As shown in Table 8 below, the F1 scores across all languages are consistently high, demonstrating the reliability of the entity extraction process. Additionally, as can be seen in Table 9, extraction accuracy is evaluated by category for English with 50 samples from each domain. Although performance varies across different domains, the overall results remain consistently strong.
@@
Italian 86.13 98.33 91.83
Japanese 89.02 97.33 92.99
Korean 88.11 96.92 92.31
Persian 83.72 98.18 90.38
Spanish 90.42 97.42 93.79
Thai 89.94 98.62 94.08
Traditional Chinese 87.15 97.50 92.04
Turkish 89.71 97.60 93.49
@@
Domain F1 (%)
Food 91.7
Beverages 91.4
Clothing 96.1
Books 97.4
Music 96.7
Transportation 97.0
Table 9: NER F1 scores across domains in English.
@@ C.2 Performance of Wikidata Search
Wikidata serves as an important source for the extraction and processing of named entities Chen et al. (2022c). We computed the proportion of extracted entities that could not be matched to any valid QID. As shown in Table 10, the average proportion of missing QIDs ranges from approximately 26% to 35% across different topics. We manually reviewed the entities that could not be matched to any valid QID and found clear signs of hallucination. In particular, the model frequently generated non-existent items, such as fictional book titles, songs, or dishes. This phenomenon poses a challenge for future efforts to manually complete missing QIDs, as hallucinated entities cannot be trivially verified or linked to real-world knowledge bases.
@@ Appendix G Results of Culture Consensus Results
In Figure 2530, we present the average consensus scores across all mentioned countries and six topics. To aid interpretation, we use yellow color coding to highlight the major Europe region. Notably, European languages consistently exhibit higher cultural consensus across models, suggesting a stronger alignment in entity representations within this region. Our findings suggest that cultural signals interact with cross-lingual representations in complex ways, shaping both specificity and consensus across languages. Approaches such as contrastive learning Chen et al. (2022a) offer a promising direction to more faithfully capturing culture-specific cues while maintaining stronger cross-lingual consensus.